F(t)= cos 5t (Laplace)


1
Expert's answer
2021-09-01T11:42:35-0400

From the table of Laplace Transforms


"L (\cos(at))=\dfrac{s}{s^2+a^2}"

Then


"L(F(t))=L (\cos(at))=\dfrac{s}{s^2+5^2}"

"L(F(t))=L (\cos(at))=\dfrac{s}{s^2+25}"